OpenStudy (anonymous):

A little car has a maximum acceleration of 2.57 m/s 2. What is the maximum acceleration of the little car if it tows another car that has the same mass?

OpenStudy (anonymous):

Your acceleration decreases by a magnitude of 2 2.57/2 =1.285N

OpenStudy (anonymous):

i believe it will still be 2.57/ a=F/m, F=f_frcition*mg. m cancels.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

if you have one car that has an acceleration of 2.57 and you double the mass by adding another car the acceleration will be half of what it is with one car

OpenStudy (jamesj):

Brent is right. The car exerts a maximum force, call it \( F_{max} \). Now as F = ma, for that max force we have \[ F_{max} = 2.57m \] Now if the mass doubles, we have \( F_{max} = (2m)a_{new} \). Because this must still be equal to the expression above, that means \[ a_{new} = \frac{2.57}{2} m/s^2 \]

OpenStudy (jamesj):

To fill in the calculation, we have \[ F_{max} = 2.57m \] and \[ F_{max} = (2m)a_{new} \] Hence \[ 2.57m = 2ma_{new} \] Cancel m from both sides and a_new is the expression above.

OpenStudy (jamesj):

@Holysmoke, welcome to OpenStudy. We're here to help, and you'll find this site most useful if you don't "drop and run": ask a question then don't stay around for the answer. It works best if you stay and engage in conversation with someone who is helping you.